OpenStudy (anonymous):

0.649 nm to mm Explain why the answer is 6.49x10^-7 mm

OpenStudy (cwrw238):

1 mm = 10^6 nm so 1 nm = 10^-6 mm 0.649 nm = 0.649 * 10^-6 nm = 6.49 * 10^-7 mm